Here's one: two cars a traveling from point A to point B and back to point A. Car A travels at 20 miles per hour from point A to point B, and 20 miles per hour from point B to point A. Car B travels 10 miles per hour from A to B, and 30 from B to A. Which car will arrive first?
Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

let's say the distance from point A to point B is 60 miles
Car A will take 60/20=3 hours each way total 6 hours
Car B will take 60/10=6 hours one way and 60/30 =2 hours to return total 8 hours
let's say the distance from point A to point B is 30 miles
Car A will take 30/20=1.5 hours each way total 3 hours
Car B will take 30/10=3 hours one way and 30/30=1 hours to return total 4 hours

Answer by Alan3354(69443)

two cars a traveling from point A to point B and back to point A. Car A travels at 20 miles per hour from point A to point B, and 20 miles per hour from point B to point A. Car B travels 10 miles per hour from A to B, and 30 from B to A. Which car will arrive first?
-------------
Avg speed of a round trip is 2*r1*r2/(r1+r2) where r1 and r2 are the speeds.
-----
For 20 & 20, it's 20 mi/hr
For 10 & 30, it's 2*10*30/(10+30) = 600/40 = 15 mi/hr
Car A will arrive first.